# CLIP-Free, Label-Free, Zero-Shot Concept Bottleneck Models

### v3

@@ Abstract
Concept Bottleneck Models (CBMs) map dense, high-dimensional feature representations into a set of human-interpretable concepts which are then combined linearly to make a prediction. However, modern CBMs rely on the CLIP model to establish a mapping from dense feature representations to textual concepts, and it remains unclear how to design CBMs for models other than CLIP. Methods that do not use CLIP instead require manual, labor intensive annotation to associate feature representations with concepts. Furthermore, all CBMs necessitate training a linear classifier to map the extracted concepts to class labels. In this work, we lift all three limitations simultaneously by proposing a method that converts any frozen visual classifier into a CBM without requiring image-concept labels (label-free), without relying on the CLIP model (CLIP-free), and by deriving the linear classifier in a zero-shot manner. Our method is formulated by aligning the original classifier’s distribution (over discrete class indices) with its corresponding vision-language counterpart distribution derived from textual class names, while preserving the classifier’s performance. The approach requires no ground-truth image–class annotations, and is highly data-efficient and preserves the classifiers reasoning process. Applied and tested on over 40 visual classifiers, our resulting CLIP-free, zero-shot CBM sets a new state of the art, surpassing even supervised CLIP-based CBMs. Finally, we also show that our method can be used for zero-shot image captioning, outperforming existing methods based on CLIP, and achieving state-of-art results.
@@ 1 Introduction
Visual classifiers predict a class as a linear combination of dense, high‑dimensional visual feature vectors that are difficult to interpret by humans. Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) address this challenge by mapping these feature vectors into a set of human-interpretable concepts, each associated with an activation score (referred to as a concept activation). Predictions are then made as a linear combination of these concept activations. Initial CBMs required image–concept annotations to train the bottleneck layer that maps dense features to concepts. Modern CBMs (oikarinen2023labelfree; Yang2022LanguageIA; Panousis2023SparseLC; Rao2024DiscoverthenNameTC) overcome this limitation by leveraging the CLIP model (Radford2021LearningTV) as an annotation-free alternative. Since CLIP models map image and text into a shared embedding space, these approaches can query image features against a set of predefined textual concepts within that space and use cosine similarity scores to find a matching annotation. These approaches are commonly referred to as label-free CBMs.
@@
Nevertheless, in many real-world scenarios a high‑performing, task‑specific legacy model already exists and typically outperforms zero‑shot CLIP models (zhang2024why). A natural question that then arises is: how to develop a label-free CBM for such legacy specialist models? Retraining such a specialist on a large image–text corpus following the CLIP approach is impractical, in terms of computational cost and need of a huge amount of image-text data. Furthermore, obtaining ground-truth image-concept annotations is time consuming and labor intensive. Finally, retraining this legacy model further alters its original decision-making process and distribution, which is typically not desired. As a result, CBMs remain confined to CLIP‑like models, leaving specialist vision models aside with no clear way to adapt them into the CBM framework.
@@
In this work, we lift this limitation by first proposing a method dubbed as TextUnlock, which aligns a distribution of a frozen visual classifier to its corresponding vision-language counterpart, without relying on CLIP. TextUnlock has four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier. The number of data points is also significantly reduced compared to CLIP-based approaches. Secondly, it is label-free, no labels are required to achieve this formulation. Thirdly, TextUnlock is trained to preserve the original distribution and reasoning process of the classifier, and does not compromise the classifier’s original performance (average of 0.2 points drop in accuracy). Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid. After the original classifier’s distribution is aligned to its vision-language counterpart distribution with TextUnlock, we simply query the transformed classifier’s image features against a set of predefined text concepts to obtain concept activations for our CBM, and then derive the concept-to-class classifier directly from the transformed classifier’s text source, without requiring any additional training but rather operating in a zero-shot manner (i.e., we do not train a linear probe to associate the concept activations to class labels).
@@
In summary, our contributions are as follows: (i) We propose a method to convert any frozen visual classifier into a CBM without relying on the CLIP bottleneck (CLIP-free), neither annotated image-concept data (label-free) while also deriving the concept-to-class classifier in a zero-shot manner. (ii) We demonstrate the effectiveness of our method with 40 different architectures, along with extensive ablation studies and intervention results, and further show how our method can also be used to perform zero-shot image captioning. (iii) Our method sets new state-of-the-art results, outperforming existing works including supervised CLIP-based CBMs, despite being trained only on ImageNet-1K.
@@ 2 Related Work
Concept Bottleneck Models (CBMs): Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) are a class of inherently interpretable models that map dense representations into human-understandable concepts, and then make predictions based on a linear combination of these interpretable units rather than opaque features. Recently, Label-Free CBMs (LF-CBMs) have been introduced, which leverage CLIP to perform the feature-to-concept mapping without requiring annotated image–concept data. This can be achieved either by using CLIP to provide ground-truth image–concept similarity scores that train the bottleneck layer (oikarinen2023labelfree), or by querying image features against a predefined set of concepts in the CLIP space (Yang2022LanguageIA; Panousis2023SparseLC) and using the resulting cosine similarities as concept activations. Several follow-up works (Rao2024DiscoverthenNameTC; Knab2024DCBMDV; Benou2025ShowAT; yuksekgonul2023posthoc) adopt this CLIP-based pipeline. However, these approaches remain limited to CLIP-based classifiers and cannot be applied to other visual models. Moreover, they require training a linear probe to map concept activations to class predictions. Our work addresses both limitations: we propose a fully CLIP-free CBM and demonstrate that our method can be readily used to derive the linear probe in a fully zero-shot manner.
Transforming Visual Features to Text: There exist works that try to decode visual features of models beyond CLIP. DeVIL (Dani2023DeViLDV) and LIMBER (merullo2023linearly) train an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data. Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process. ZS-A2T (Salewski2023ZeroshotTO) converts attention maps into natural language in a zero-shot manner using Large Language Models (LLMs), but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective). Text-to-Concept (T2C) (Moayeri2023TextToConceptB) also trains a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder. Notably, all these methods rely on the CLIP approach and/or its supervision, and they alter the classifier’s distribution by entirely discarding its output class distribution. In contrast, our method is CLIP-free, can be applied to any pretrained classifier without requiring any annotated data, and explicitly preserves the classifier’s predictive distribution.
@@ 3 Method
We first elaborate on our proposed TextUnlock method in Section 3.1, which will then allow us to design zero-shot concept bottleneck models, which we discuss in Section 3.2.
@@ 3.1 TextUnlock
A visual classifier assigns an image to a specific category from a predefined set of discrete class labels. For example, in ImageNet-1K (Deng2009ImageNetAL), this set contains $1,000$ class labels. Originally, these discrete class labels correspond to class names in text format. For example, in ImageNet-trained models, the discrete label $1$ corresponds to the class goldfish. These classes are typically discretized to facilitate training with cross-entropy. However, when the textual class names are used, they provide an advantage. Specifically, when the textual class names are embedded into vector representations (e.g., using a word embedding model or text encoder), they provide semantic information. Specifically, These embeddings reside in a continuous space where nearby vectors capture related semantic associations. For the “goldfish” example, neighboring vectors might include terms such as “freshwater”, “fins” and “orange”. Such associations can be viewed as high-level conceptual attributes that characterize the class. Our method learns to map images into this text embedding space using only class names, thus linking both the class name and its surrounding semantic associations with the image. This process thus naturally supports unseen words that are not part of the class names (e.g., concepts in the CBM). We accomplish this through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to match its distribution with the original classifier’s class distribution. This is done while keeping both the visual and textual encoders frozen. By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space. This allows us to query the visual classifier with text queries beyond the class names (e.g., concepts) and obtain concept activations for CBMs.
@@
Figure 1: Overview of our proposed TextUnlock. (a) The process of training the MLP mapping between vision and text space with pseudocode given in Appendix Section A. (b) The process of inference with the adapted visual classifier. The text encoder acts as weight generator for a linear classifier. Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
@@
The most straightforward way to training the MLP is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution. However, this approach violates two key desiderata: (1) it necessitates annotated data, and (2) re-training the legacy classifier alters its original decision distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions). Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored. We instead propose to align $S$ to the original decision distribution $o$ through cross-entropy loss. For a single sample, the loss is given by
@@
This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but instead distill the distribution of the original model to its counterpart vision-language distribution. Note that the loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters. Eq. 1 shows that our approach is label-free and steers the MLP to be faithful to the original classifier $F$, since it is explicitly trained for that purpose. Note also that this loss function naturally encodes the classifier’s relationship across all classes.

### v4

# CLIP-Free, Label Free, Unsupervised Concept Bottleneck Models

@@ Abstract
Concept Bottleneck Models (CBMs) map dense feature representations into human-interpretable concepts which are then combined linearly to make a prediction. However, modern CBMs rely on the CLIP model to obtain image-concept annotations, and it remains unclear how to design CBMs without the CLIP bottleneck. Methods that do not use CLIP instead require manual, labor intensive annotation to associate feature representations with concepts. Furthermore, all CBMs necessitate training a linear classifier to map the extracted concepts to class labels. In this work, we lift all three limitations simultaneously by proposing a method that converts any frozen visual classifier into a CBM without requiring image-concept labels (label-free), without relying on the CLIP model (CLIP-free), and by deriving the linear classifier in an unsupervised manner. Our method is formulated by aligning the original classifier’s distribution (over discrete class indices) with its corresponding vision-language counterpart distribution derived from textual class names, while preserving the classifier’s performance. The approach requires no ground-truth image–class annotations, and is highly data-efficient and preserves the classifiers reasoning process. Applied and tested on over 40 visual classifiers, our resulting unsupervised, label-free and CLIP-free CBM (U-F2-CBM) sets a new state of the art, surpassing even supervised CLIP-based CBMs. We also show that our method can be used for zero-shot image captioning, outperforming existing methods based on CLIP, and achieving state-of-art.
@@ 1 Introduction
Visual classifiers predict a class as a linear combination of dense, high‑dimensional visual feature vectors that are difficult to interpret by humans. Concept Bottleneck Models (CBMs) [20] address this challenge by mapping these feature vectors into a set of human-interpretable concepts, each associated with an activation score (referred to as a concept activation). Predictions are then made as a linear combination of these concept activations. Initial CBMs required image–concept annotations to train the bottleneck layer that maps dense features to concepts. Modern CBMs [36, 60, 5] overcome this limitation by leveraging the CLIP model [43] or vision–language grounding models [48] to provide image-concept annotations. Since CLIP models map image and text into a shared embedding space, these approaches can query images against an entire pool of predefined textual concepts and use cosine similarity scores to find a matching annotation. These approaches are commonly referred to as label-free CBMs.
@@
In many real-world settings, a high-performing, task-specific legacy model already exists and often achieves strong results on the target task [63]. When CLIP is used to generate image-concept annotations, the resulting CBM is anchored to CLIP’s embedding space; the legacy model must be interpreted through CLIP’s notion of similarity, rather than through its own learned representation. In this sense, the approach is not fundamentally different from replacing the legacy model with a CLIP-based CBM. Furthermore, this dependence can transfer CLIP’s biases and behavior into the legacy model (e.g., importing the CLIP typographic bias [11] into a DINO model). A natural question that then arises is: how to develop a label-free CBM for such legacy specialist models without the CLIP constraint? Retraining such a specialist on a large image–text corpus following the CLIP approach is impractical, in terms of computational cost and need of a huge amount of image-text data. Furthermore, obtaining ground-truth image-concept annotations through manual human-labor is time consuming and expensive. Finally, retraining this legacy model further alters its original decision-making process and distribution, which is typically not desired.
@@
In this work, we lift these limitations by first proposing a method dubbed as TextUnlock, which aligns a distribution of a frozen visual classifier to its corresponding vision-language counterpart, without relying on CLIP. TextUnlock has four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier. The number of data points is also significantly reduced compared to CLIP-based approaches. Secondly, it is label-free, no labels are required to achieve this formulation. Thirdly, TextUnlock is trained to preserve the original distribution and reasoning process of the classifier, and does not compromise the classifier’s original performance (average of 0.2 points drop in accuracy). Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid. After the original classifier’s distribution is aligned to its vision-language counterpart distribution with TextUnlock, we simply query the transformed classifier’s image features against a set of predefined text concepts to obtain concept activations for our CBM, and then derive the concept-to-class classifier directly from the transformed classifier’s text source, without requiring any additional training and operating in an unsupervised manner (i.e., we do not train a linear probe to associate the concept activations to class labels).
Figure 1: Overview of our proposed TextUnlock. (a) The process of training the MLP mapping between vision and text space. (b) The process of inference with the adapted visual classifier. The text encoder acts as weight generator for a linear classifier. Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
@@
In summary, our contributions are as follows: (i) We propose a method to convert any frozen visual classifier into a CBM without relying on the CLIP constraint (CLIP-free), neither on annotated image-concept data (label-free), and for the first time, also deriving the concept-to-class classifier in a fully unsupervised manner. (ii) We demonstrate the effectiveness of our method with 40 different architectures, along with extensive ablation studies and intervention results, and further show how our method can also be used to perform zero-shot image captioning. (iii) We set new state-of-the-art results on CBMs, outperforming existing works including supervised CLIP-based CBMs, despite being trained only on ImageNet-1K.
@@ 2 Related Work
Concept Bottleneck Models (CBMs) are inherently interpretable models that predict by first mapping inputs to human-understandable concepts and then combining those concepts linearly to make the final decision, originally requiring concept annotations for each image [36]. Recently, Label-Free CBMs (LF-CBMs) have been introduced, which leverage CLIP to provide image-concept annotations. CLIP is used to generate ground-truth image–concept similarity scores, which serve as supervision for training the bottleneck layer of a downstream classifier. Several further works [60, 5] adopted this pipeline. Another work uses vision-language grounding models [48] to obtain image-concept annotations. Another line of work [59, 39, ao2024DiscoverthenNameTC, 19] builds CBMs directly on top of CLIP, making them specifically tailored to CLIP-only models and thereby eliminating the need for a separate bottleneck training stage. In these methods, image features are queried against a predefined set of concepts in the CLIP embedding space, and the resulting cosine similarities are used directly as concept activations. As a result, all the above methods either (1) are limited to CLIP models, or (2) depend on CLIP or vision–language models to obtain the image-concept annotations. On the other hand, we propose a fully CLIP-free CBM; our method does not rely on any external vision-language model to generate the image-concept annotations, nor on human annotation, and can be applied to any legacy model without the CLIP constraint. Finally, all CBMs works to date require training a linear probe to map concept activations to class predictions. In contrast, we demonstrate that our method can be readily used to derive the linear probe in a fully unsupervised manner.
Transforming Visual Features to Text: There exist works that try to decode visual features of models beyond CLIP. DeVIL [7] and LIMBER [34] train an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data. Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process. Text-to-Concept (T2C) [35] also trains a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder. Notably, all these methods rely on the CLIP approach and/or its supervision, and they alter the classifier’s distribution by entirely discarding its output class distribution. In contrast, our method is CLIP-free, can be applied to any pretrained classifier without requiring any annotated data, and explicitly preserves the classifier’s predictive distribution.
@@ 3 Method
We first elaborate on our proposed TextUnlock method in Section 3.1, which will then allow us to design our proposed U-F2-CBM, which we discuss in Section 3.2. U stands for unsupervised; L2 for CLIP-free and Label-free (hence ‘double free’).
@@ 3.1 TextUnlock
A visual classifier assigns an image to a specific category from a predefined set of discrete class labels. For example, in ImageNet-1K [9], this set contains $1,000$ class labels. Originally, these discrete class labels correspond to class names in text format. For example, in ImageNet-trained models, the discrete label $1$ corresponds to the class goldfish. These classes are typically discretized to facilitate training with cross-entropy. However, when the textual class names are used, they provide an advantage. Specifically, when the textual class names are embedded into vector representations (e.g., using a word embedding model or text encoder), they provide semantic information. Specifically, These embeddings reside in a continuous space where nearby vectors capture related semantic associations. For the “goldfish” example, neighboring vectors might include terms such as “freshwater”, “fins” and “orange”. Such associations can be viewed as high-level conceptual attributes that characterize the class. Our method learns to map images into this text embedding space using only class names, thus linking both the class name and its surrounding semantic associations with the image. This process thus naturally supports unseen words that are not part of the class names (e.g., concepts in the CBM). We accomplish this through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to match its distribution with the original classifier’s class distribution. This is done while keeping both the visual and textual encoders frozen. By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space. This allows us to query the visual classifier with text queries beyond the class names (e.g., concepts) and obtain concept activations for CBMs.
@@
Figure 2: For any pretrained classifier, we first perform (a) concept discovery, followed by (b) building the concepts-to-class classifier in an unsupervised manner, which results in (c) our final U-F2-CBM. Note that the concept bank only needs to be encoded once.
The most straightforward way to training the MLP is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution. However, this approach violates two key desiderata: (1) it necessitates annotated data, and (2) re-training the legacy classifier alters its original decision distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions). Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored. We instead propose to align $S$ to the original decision distribution $o$ through cross-entropy loss. For a single sample, the loss is
@@
This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but instead distill the distribution of the original model to its counterpart vision-language distribution. This loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters. Eq. 1 shows that our approach does not require any labels for training, and steers the MLP to preserve the classifier’s distribution, since it is explicitly trained for that purpose. Note also that this loss function naturally encodes the classifier’s relationship across all classes.